The first tern of an A.P. is 2 and the 6th term is 12. Find the

sum of the first 12 terms.


1
Expert's answer
2021-09-17T03:50:06-0400

a1+(n1)d.....Apseriesa1 +(n-1)d.....Ap series

sumsum

sn=0.5n(2a+(n1)d)sn = 0.5n(2a+(n-1)d)

a1=2a1 = 2

a1+5d=12a1 + 5d = 12

d=2d= 2

0.512(22+(121)2)0.5*12(2*2 + (12-1)2)

sum=156sum = 156